### v1

# Preference-driven Knowledge Distillation for Few-shot Node Classification

Abstract
Graph neural networks (GNNs) can efficiently process text-attributed graphs (TAGs) due to their message-passing mechanisms, but their training heavily relies on the human-annotated labels. Moreover, the complex and diverse local topologies of nodes of real-world TAGs make it challenging for a single mechanism to handle. Large language models (LLMs) perform well in zero-/few-shot learning on TAGs but suffer from a scalability challenge. Therefore, we propose a preference-driven knowledge distillation (PKD) framework to synergize the complementary strengths of LLMs and various GNNs for few-shot node classification. Specifically, we develop a GNN-preference-driven node selector that effectively promotes prediction distillation from LLMs to teacher GNNs. To further tackle nodes’ intricate local topologies, we develop a node-preference-driven GNN selector that identifies the most suitable teacher GNN for each node, thereby facilitating tailored knowledge distillation from teacher GNNs to the student GNN. Extensive experiments validate the efficacy of our proposed framework in few-shot node classification on real-world TAGs. Our code is available at https://github.com/GEEX-Weixing/PKD.
1 Introduction
Text-attributed graphs (TAGs yang2021graphformers ), such as citation, webpage, and product graphs feng2024taglas ; li2025hetgb , have nodes associated with text attributes. Graph neural networks (GNNs) kipf2016semi ; yun2019graph have demonstrated excellent performance and efficiency in node classification on TAGs, which are supported by high-quality labels and effective message-passing mechanisms zhu2021textgnn . However, the manual labeling of nodes is undoubtedly a tedious, expensive, and time-consuming task shi2025latent ; cao2025low . In many scenarios, only a few node labels are available. Additionally, nodes often have complex and diverse interaction relationships with each other—their local topologies are intricate—which challenge traditional GNNs with fixed message-passing mechanisms. Compared with GNNs, large language models (LLMs) exhibit impressive zero-/few-shot learning capabilities on TAGs chen2024exploring ; wang2024llms ; wu2024exploring . But the large parameter scale considerably hinders their inference efficiency he2023harnessing .
A natural idea is to blend their complementary strengths for few-shot node classification on TAGs. Knowledge distillation (KD) hinton2014distilling is a feasible solution. However, directly distilling knowledge from the LLM to GNN is impractical. Firstly, the discrepancy of decoder-only (LLMs) and encoder-only (GNNs) leads to fundamentally different characteristics in their embedding spaces chen2023label . And the huge embedding-dimension difference needs sophisticated embedding alignment and also brings high training cost hu2025large . In contrast, conducting prediction distillation from LLMs to GNNs by annotating node labels can efficiently alleviate the label scarcity and scalability dilemma pan2024distilling . The critical question is how to select the nodes for the LLM’s label annotation to effectively enhance teacher GNNs. Generally, one may use uncertainty fox1957training as a selection metric in the embedding space of GNN. However, owing to nodes’ diverse semantic and complex structural attributes (e.g., local topologies), a single GNN cannot capture the essences of nodes completely liu5084903distilling . Therefore, we investigate the embedding spaces of various-architecture GNNs to effectively mitigate cognitive limitations wang2021mulde associated with relying on a single GNN, thereby better selecting nodes for LLM’s label annotations.
Nevertheless, since nodes have intricate local topologies, which need tailored message-passing mechanisms, how to tailor for each node the most appropriate message-passing mechanism is another challenge. Different GNNs provide different prediction attributes for each node during the learning process jin2023multi , encompassing the understandings of its topologies, its interaction relationships to other nodes, and its latent patterns. These node-specific attribute differences suggest that a single message-passing mechanism cannot fundamentally handle the entire graph. Some studies guo2023boosting ; liu5084903distilling distill knowledge sequentially or simultaneously from teacher GNNs without taking into account the node-specific local topologies, resulting in no obvious performance improvement or even performance degradation yang2024two . Therefore, it is essential to identify the GNN message-passing mechanisms that align with the node-specific attributes.
To this end, we propose a preference-driven knowledge distillation (PKD) framework that unites the complementary strengths of LLMs and various-architecture GNNs for few-shot node classification on TAGs. It mainly includes two modules: GNN-preference-driven Node Selector (GNS) and Node-preference-driven GNN Selector (NGS). The prerequisite of GNS is that the LLM should be able to comprehend the graph topology. Thus, we develop the graph topology aware (GTA) prompts to fine-tune the LLM, enhancing its capacity to comprehend graph topology. GNS fully exploits nodes’ prediction discrepancies among various GNNs to decide nodes whose labels are annotated by the LLM will effectively enhance teacher GNNs, facilitating knowledge distillation from the LLM to teacher GNNs. NGS selects for each node the most appropriate GNN message-passing mechanism, facilitating the tailored knowledge distillation from various teacher GNNs to the student GNN. It regards the fine-tuned LLM as the RL-based (reinforcement learning) agent, which treats all textualized node-specific attributes (including node’s semantic, structure, and prediction attributes) as state and the student GNN’s performance as reward.
Contributions
❶We introduce a preference-driven knowledge distillation (PKD) framework to synergize the complementary strengths of the LLM and various GNNs ingeniously for few-shot node classification on TAGs; ❷ we propose a GNN-preference-driven node selector, effectively determining nodes for annotation by the LLM and promoting knowledge distillation from the LLM to teacher GNNs; ❸ we propose a node-preference-driven GNN selector to tailor for each node the most appropriate message-passing mechanism, promoting knowledge distillation from teacher GNNs to the student GNN; ❹ we validate the efficacy of PKD for few-shot node classification on nine TAGs. The experiments show that it even defeats some state-of-the-art methods that use more node labels.
2 Related Work
2.1 Graph Neural Networks
The field of graph learning has been dominated by GNNs, such as GCN kipf2016semi , GAT velivckovic2017graph , APPNP gasteiger2018predict , $\text{H}_{2}\text{GCN}$ zhu2020beyond , and GPRGNN chien2020adaptive . Recently, HoloNets koke2023holonets introduces a dual-filter mechanism with spectral response, extending spectral convolutions to directed graphs. DirGNN rossi2024edge defines the in-neighbors and out-neighbors and performs separate propagation and aggregation, improving the message passing through the incorporation of edge directionality. To deal with label scarcity, GCNII chen2020simple introduces initial residual connections and identity mapping to construct a deep GNN while EGNN satorras2021n enforces equivariance constraints for the enhancement of data efficiency and generalization. AGST ding2024toward and IceBerg li2025iceberg leverage the different self-training li2018deeper methods to effectively utilize unlabeled nodes.
2.2 Knowledge Distillation
KD is not only used for model compression, but for strengthening purposeful abilities of the student model. GFL yao2020graph extracts structural knowledge from a pre-prepared similar auxiliary graph, distilling it to the target graph for enhancing few-shot node classification performance. KDGA wu2022knowledge utilizes multiple graph augmentation strategies to make student GNN produce robust node representations after distillation. MSKD zhang2022multi mitigates the diverse classification situations requiring for different nodes by capturing multi-scale topological semantics distilled from varying layers. However, the capability of an individual teacher is inherently limited. BGNN guo2023boosting distills complementary knowledge from multiple GNN teachers sequentially and integrate it by the adaptive temperature parameter and weight boosting modules. MTAAM yang2024two distills knowledge of multiple teacher GNNs into an MLP-student, offering quick inference speed without compromising accuracy. FairGKD zhu2024devil obtains equitable and informative node representations by synergizing multiple GNN experts into a teacher. DMKD liu5084903distilling harnesses complementary knowledge from various GNNs and conducts layer-level knowledge distillation to mitigate the constraint of a single teacher. Furthermore, chen2023label is a label-free method that proposes the LLM-GNN. It uses LLMs to get high-quality annotation through active and confidence-awareness node selection, thereby circumventing the difficulty of label annotation by humans. LinguGKD hu2025large introduces a kind of ingenious contrastive learning to align the LLM’s semantic features with GNN’s structural features to achieve knowledge transfer. Most of the above knowledge distillation methods do not tailor for each node the most appropriate message-passing mechanism and underperform on few-shot node classification.
3 Method: PKD
In this section, we present the preference-driven knowledge distillation (PKD) framework. PKD involves two key modules: GNN-preference-driven Node Selector (GNS) and Node-preference-driven GNN Selector (NGS). The main goal of the former module is to select node groups whose labels are annotated by the LLM will drastically enhance teacher GNNs. The main goal of the latter module is to select the most appropriate teacher GNN for each node, thereby tackling the complication of node-specific local topologies. The PKD framework is illustrated in detail in Figure 1.
Figure 1: Overview of PKD. The framework has two key modules: GNN-preference-driven Node Selector (GNS) and Node-preference-driven GNN Selector (NGS). Before starting GNS, we first fine-tune the LLM with GTA prompts to enable it to comprehend graph properties. In the GNS module, we exploit the proposed $K$-uncertainty based on the node prediction uncertainty in each teacher GNN’s embedding space to select nodes. For effectively exploiting the LLM to annotate those selected nodes, we combine the semantic attributes and structure attributes derived from the proposed Distance-based Neighbor Selector (DNS) module on these nodes to construct prompt, promoting the prediction distillation from the fine-tuned LLM to teacher GNNs ($T_{1},T_{2},\ldots,T_{B}$). In the NGS module, we select for each node the most appropriate teacher GNN for tailored knowledge distillation. The teacher GNN selection is achieved by reinforcement learning with the fine-tuned LLM as agent.
3.1 Background
A text-attributed graph (TAG) is denoted by $\mathcal{G}_{T}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{A},\mathbf{T})$, where $\mathcal{V}=\left\{v_{1},\ldots,v_{N}\right\}$ is a set of nodes with semantic attributes $\mathbf{T}=\left\{\mathbf{t}_{1},\ldots,\mathbf{t}_{N}\right\}$ and $\mathcal{E}$ is a set of edges. Each semantic attribute can then be encoded as a sentence embedding $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{i},\ldots,\mathbf{x}_{N}]\in\mathbb{R}^{N\times F}$ with the help of language models. $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix. Given the few-shot node classification task, let $\mathcal{D}_{L}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{Q}\;(Q\ll N)$ be the set of labeled nodes with $\mathbf{y}_{i}$ as the one-hot label of the training sample $\mathbf{x}_{i}$ and $\mathcal{D}_{U}$ be the set of unlabeled nodes, respectively. The goal is to accurately predict the labels of nodes that belong to $\mathcal{D}_{U}$ given few labeled nodes in $\mathcal{D}_{L}$.
We assume $B$ teacher GNNs denoted by $\{T_{b}\}_{b=1}^{B}$, and $f_{T_{b}}^{\theta}$ is the model parameters of $T_{b}$. The $B$ logit outputs of teacher GNNs for node $v_{i}$ are written as $\mathbf{z}_{i}^{T}=[\mathbf{z}_{i,1}^{T},\ldots,\mathbf{z}_{i,b}^{T},\ldots,\mathbf{z}_{i,B}^{T}]$, which is the concatenation of the logit of each teacher $\mathbf{z}_{i,b}^{T}=[z_{i,b,1}^{T},\ldots,z_{i,b,c}^{T},\ldots,z_{i,b,C}^{T}]\;(1\leq b\leq B)$, where $z_{i,b,c}^{T}$ is the probability of $v_{i}$ belonging to class $c\;(1\leq c\leq C)$ computed by teacher $T_{b}$. Our final objective for the KD from node-preference GNNs to the student GNN can be divided into three parts:
$\mathcal{L}_{KD}=\alpha\cdot(-\frac{1}{N}\sum_{i}^{N}\widetilde{\mathbf{z}}_{i}^{T}\cdot\text{log}f_{S}^{\theta}(\mathbf{x}_{i}))+\beta\cdot(-\frac{1}{Q}\sum_{i}^{Q}\mathbf{y}_{i}\cdot\text{log}f_{S}^{\theta}(\mathbf{x}_{i}))+\gamma\cdot(\frac{1}{N}\sum_{i=1}^{N}H(f_{S}^{\theta}(\mathbf{x}_{i})))$ (1)
where $\alpha,\beta,\gamma$ are hyper-parameters to balance three losses. For student GNN with parameters $f_{S}^{\theta}$, the first loss, distillation loss $\mathcal{L}_{DL}$, is defined as the cross-entropy between the predictions of the teacher GNNs and that of the student GNN. The $f_{S}^{\theta}(\mathbf{x}_{i})$ is the Softmax output of student GNN and it denotes the probability distribution of $v_{i}$ belonging to class $c$. $\widetilde{\mathbf{z}}_{i}^{T}=\mathbf{m}_{i}\otimes\mathbf{z}_{i}^{T}$, where $\mathbf{m}_{i}$ is a one-hot vector denoting which teacher GNN is preferred by $v_{i}$. The second loss, $\mathcal{L}_{CE}$, is the cross-entropy loss in the training of student GNN. Inspired by ying2018hierarchical , we add $\mathcal{L}_{E}$ to the objective as the last part, which makes the logits of student GNN closer to one-hot vectors. The $H(\cdot)$ denotes Shannon entropy.
Figure 2: The performance improvements in zero-shot node classification on homophily and heterophily graphs.
3.2 LLM Fine-tuning
Recent studies reveal that LLMs possess reasoning apabilities wang2024can , but they often underperform compared to even the simple GNNs when tackling graph learning. The key challenge lies in its inability to directly process the raw graph data and understand topology properties, limiting the generalization ability of LLMs in this domain. To address this, we propose GTA prompts fine-tuning.
This method consists of four distinct fine-tuning instruction types, each designed to enhance structural comprehension, such as local connectivity, node degree, cycle structure, and path-based dependencies, by addressing specific tasks: (1) Connectivity involves determining whether or not two nodes in an undirected graph are connected; (2) Degree requires the LLM to determine the degree of a given node based on the adjacency matrix $\mathbf{A}$; (3) Cycle Detection requires the LLM to ascertain whether a cycle exists within the given sequence of nodes; (4) Text Generation demands the LLM to generate textual contents of given nodes based on the semantic attributes of preceding nodes in the random walk. Through fine-tuning, LLM exhibits significant improvements on the zero-shot node classification task, as demonstrated in Figure 2. More detailed task descriptions and detailed task-specific GTA prompt templates are provided in Appendix B.3.
3.3 GNN-preference-driven Node Selector
After being fine-tuned, the LLM can generate superior node label annotations (as shown in Figure 2). However, how to select nodes for LLM’s label annotation to effectively enhance teacher GNNs (those nodes are assumed to be preferred by GNNs) is a challenging problem. Uncertainty is an essential metric for node selection. It mainly consists of two parts: random uncertainty caused by inherent noise and cognitive uncertainty caused by insufficient observation. The former type is inevitable, so we focus on the latter type. From the perspective of collective consensus fan2023collaborative , we design the GNN-preference-driven Node Selector based on the defined $K$-uncertainty ($\delta_{K}$). Specifically, we measure the cognitive disagreement among the teacher GNNs’ SoftMax outputs using the Kullback-Leibler (KL) divergence, and get the preference ranks of all nodes by $\delta_{K}$, i.e., $\mathcal{V}_{\mathcal{PR}}=\text{Sort}({\left\{v_{1},\ldots,v_{N}\right\},\delta_{K}(v_{1}),\delta_{K}(v_{2}),\ldots,\delta_{K}(v_{N})})$. High $K$-uncertainty of nodes indicates that their prediction uncertainty by GNNs is higher. Those nodes can effectively enhance GNNs if their more accurate labels, annotated by the LLM, are provided to train GNNs, as the following proposition suggests.
Proposition 3.1.
These nodes with higher $K$-uncertainty ($\delta_{K}$) are beneficial for GNNs enhancement.
$\delta_{K}(v)\triangleq\sum_{1\leq i<j\leq B}^{B}[D_{KL}(f_{T_{i}}^{\theta}(v)||f_{T_{j}}^{\theta}(v))+D_{KL}(f_{T_{j}}^{\theta}(v)||f_{T_{i}}^{\theta}(v))]\varpropto\delta_{v}$ (2)
where $\delta_{v}$ is the uncertainty of node $v$, is defined as $\frac{1}{B}\sum_{i=1}^{B}D_{KL}(f_{T_{i}}^{\theta}(v)||\mathcal{M}(v))$. The $\mathcal{M}(v)$ is the average prediction probability distribution of all $B$ teacher GNNs (See Definition 11 for details). $D_{KL}(\cdot||\cdot)$ is the function to calculate KL divergence.
${f_{T}^{\theta}}^{*}(\tilde{\mathcal{D}}_{L})=\mathop{\arg\min}\limits_{v_{i}\in\{v_{\mathcal{PR}}^{1},v_{\mathcal{PR}}^{2},\ldots,v_{\mathcal{PR}}^{W}|\delta_{K}(v_{\mathcal{PR}}^{W})>\tilde{\delta}_{K}\}}\frac{1}{W}\sum\mathcal{L}(f_{T}^{\theta},v_{i})$ (3)
where $\tilde{\mathcal{D}}_{L}$ is the expanded training dataset. $f_{T}^{\theta}$$^{*}$ is the optimal parameter of teacher GNN. $v_{\mathcal{PR}}^{w}$ represents the $w$-th nodes in the preference rank. $W$ is the number of selected nodes by GNS and the $\tilde{\delta}_{K}$ is the $K$-uncertainty threshold depending on the expansion ration.
Figure 3: This is exemplified using the Cora dataset. Starting from the arrow and progressing counterclockwise, the KL divergence sum gradually increases, accompanied by a darkening of the triangle colors. The length of each triangle indicates the number of nodes within a specific KL divergence sum range, where $N_{P}$ denotes the number of classes predicted by the teacher GNNs.
The proof is given in Appendix D. By selecting these nodes (illustrated in Figure 3), we ensure that the most uncertain and informative nodes are labeled by the LLM to promote the progress of prediction distillation through the cross-entropy function. Correspondingly, GNS also reduces the inference costs associated with LLMs by not querying all nodes in $\mathcal{D}_{U}$. To generate high-quality annotations for GNN-preferred nodes, we further design the Distance-based Neighbors Selection (DNS) module, which performs the K-Nearest Neighbor (KNN) search around each selected node across the embedding spaces generated by pretrained teacher GNNs and deletes repeated neighbors. The structure attributes composed of selected neighbors and their textual contents are integrated into the category-induction prompt and inputted into the LLM. Unlike relying solely on neighbors identified by the adjacency matrix (prone to biases from 1-hop homophily), our approach ensures a more robust and diverse selection of high-quality neighbors, facilitating better construction of the category-induction prompt for the LLM. We do not select common KNN neighbors across all the embedding spaces generated by the teacher GNNs, as they may overfit to the adjacency structure.
3.4 Node-preference-driven GNN Selector
Distilling knowledge simultaneously from multiple teachers to the student is not a good option since nodes with varying local topologies require distinct message-passing modes for optimal representation updates. To achieve this, we introduce the Node-preference-driven GNN Selector (NGS) to select the most appropriate teacher for each node according to the specific attributes and promote tailored knowledge distillation. For each node in the expanded training data (including the initial few labeled nodes and those selected nodes whose labels are annotated by the LLM), we construct a node-specific prompt by combining its semantic, structural, and prediction attributes derived from the enhanced teacher GNNs. This prompt is then inputted to the fine-tuned LLM to determine the most suitable teacher for this node. The GNN selection task is formulated as a reinforcement learning problem that needs to explore the discrete action space and find a series of assignment actions to get the highest global reward across the expanded training data. Through interaction with the training process, the selector progressively refines its decisions on node-to-teacher assignments, leading to a more efficient and effective assignment strategy. Specifically, the fine-tuned LLM, serving as the agent, selects the most appropriate teacher for each node. The policy is trained to maximize classification accuracy on the expanded training data, with the reward tied to the student’s performance. To address the non-differentiability of the LLM’s decoding process, we add two additional projectors (MLPs) after the logit layer to generate action probabilities and corresponding value estimations, enabling the agent to take discrete teacher-selection actions.
In the RL framework, the elements are structured as (State, Action, Reward). During each iteration, the agent interacts with the environment by receiving all attributes of one node in the expanded training data. The agent then takes an action on which teacher is more appropriate.
State: Each state corresponds to the prompt $\mathcal{P}_{i}$ of a node, including node-specific semantic, structural, and prediction attributes. These prompts are detailed in Appendix B.2. The size of the expanded training data is denoted as $W$.
Action: The Policy Model (the fine-tuned LLM combined with an MLP projector) generates a text-related output to indicate its selection from multiple teachers, formulated as a probability distribution vector $\bm{\pi}_{T}=[\pi_{T_{1}},\pi_{T_{2}},\ldots,\pi_{T_{B}}]$, where $\pi_{T_{b}}$ denotes the probability of selecting the $b$-th teacher $T_{b}$. The action is determined through sampling.
Reward: The function is correlated with the performance of the student GNN, which is trained by distilling knowledge from the selected teacher for each node. The reward function consists of three key parts: classification accuracy, cross-entropy loss, and distillation loss. It can be written as follows:
$R=\eta\ast(\mathcal{L}_{DL}^{{}^{\prime}}-\mathcal{L}_{CE})+(1-\eta)\ast A_{cc}$ (4)
where $A_{cc}$ represents the classification accuracy of the student GNN on the expanded training data, $\eta$ is a hyper-parameter to balance the three parts, where $\mathcal{L}_{DL}^{{}^{\prime}}=-\frac{1}{W}\sum_{i}^{W}\widetilde{\mathbf{z}}_{i}^{T}\cdot\text{log}f_{S}^{\theta}(\mathbf{x}_{i})$, and $\mathcal{L}_{CE}=-\frac{1}{Q}\sum_{i}^{Q}\mathbf{y}_{i}\cdot\text{log}f_{S}^{\theta}(\mathbf{x}_{i})$.
To effectively optimize the agent’s actions for better knowledge distillation, we employ the simplified version of Proximal Policy Optimization (PPO) schulman2017proximal algorithm, which retains the core principles. Specifically, we do not instantiate the Reward Model explicitly and calculate the reward based on the performance of the student GNN. The Reference Model is also not explicitly referenced, because the parameter update objective function we utilize involves a comparison with the previous strategy. To avoid large fluctuations between the current and old policies, we adopt the CLIP strategy schulman2017proximal to limit the update margin. During the KD process, the parameters $f_{\mathcal{A}}^{\theta}$ of NGS, remain fixed, while the parameters $f_{S}^{\theta}$ of the student GNN are trained. During the NGS process, the parameters $f_{S}^{\theta}$ of the student GNN are kept fixed to compute the reward, while the parameters $f_{A}^{\theta}$ of NGS based on the collected rewards from all episodes are optimized. The pseudocode, detailed implementations, and time complexity analysis are provided in Appendix C.
4 Experiments
4.1 Experimental Setup
Datasets
In order to assess the few-shot node classification performance of our method on TAGs, we conduct a comprehensive series of experiments across 9 real-world datasets: Cornell, Washington, Texas, Wisconsin zhu2020beyond , Amazon Ratings platonov2023critical , Ogbn-Arxiv hu2020open , Wiki CS mernyei2020wiki , Pubmed, Cora yang2016revisiting . They have various 1-hop homophily ratios yang2022incorporating and additional details of the datasets can be found in Appendix A. For the KD-baselines, we partition the nodes of each graph into training, validation, and test sets, allocating 48%, 32%, and 20%, respectively, based on the proportion division mentioned in liang2024sign . For PKD and other baselines, we randomly select 1, 3, and 5 labeled nodes per class as the initial training data and then expand the dataset to 48% of the total using the GNS module. The remaining data is randomly split into 32% for validation and 20% for testing, with the preserved indices for the baselines. This operation is repeated 5 times. We report the average test classification accuracy and standard deviation of each model with parameters that lead to the peak validation accuracy.
Baselines
We compare our method against the following baseline models: (i) Advanced GNNs: GCNII chen2020simple and EGNN satorras2021n ; (ii) GNNs enhanced by LLMs: LLMGNN chen2023label and GAugLLM fang2024gaugllm ; (iii) self-training for graph learning: Self-training li2018deeper , AGST ding2024toward and IceBerg li2025iceberg ; (iv) Knowledge Distillation (KD) for GNNs: KDGA wu2022knowledge , MSKD zhang2022multi , BGNN guo2023boosting , MTAAM yang2024two , and FairGKD zhu2024devil . For homophily graphs, the teacher GNNs used are: GCN kipf2016semi ($T_{1}$), GAT velivckovic2017graph ($T_{2}$), APPNP gasteiger2018predict ($T_{3}$), $\text{H}_{2}\text{GCN}$ zhu2020beyond ($T_{4}$), and the student is GCN; for heterophily graphs, the teacher GNNs employed are: DirGNN rossi2024edge ($T_{1}$), GPRGNN chien2020adaptive ($T_{2}$), HoloNets koke2023holonets ($T_{3}$), $\text{H}_{2}\text{GCN}$ ($T_{4}$) and the student is $\text{H}_{2}\text{GCN}$. The LLM used in the experiments is Llama-3.1-8B-Instruct dubey2024llama .
4.2 Performance Analysis and Discussion
Notably, # LN 1, # LN 3, # LN 5 indicate only 1, 3, 5 labeled nodes per class are used for training PKD, while the results of the teacher GNNs ($\{T_{i}\}_{i=1}^{4}$) and other baselines are trained under the data splitting of 48%/32%/20% as mentioned above. According to Table 1, our method almost achieves the best or second-best accuracy results.
Table 1: Node classification accuracies (%) on real-world datasets. $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$ denote the teacher GNNs for homophily or heterophily graphs (refer to the descriptions in Baselines for more details of the teacher and student GNNs). The OOM stands for Out-Of-Memory. The best results are highlighted in dark gray, while the runner-up results are marked in light gray.
Methods Dataset Cornell Washington Texas Wisconsin
Amazon
Ratings
Ogbn-
Arxiv
Wiki CS Pubmed Cora
[…]
Due to the extreme insufficiency of labels, GCNII and EGNN are restricted in further improvement, although they have distinctive network architectures. Lacking carefully designed fine-tuning and enough cognition makes LLMGNN fail to produce high-quality pseudo labels and is dramatically defeated by our method PKD. Although GAugLLM harnesses LLM for feature and structure augmentations to benefit GNN, its self-training depends only on SoftMax scores to identify candidate nodes to assign pseudo-labels, a method that can sometimes be unreliable. GAugLLM achieves the best result on the Pubmed dataset, but it is outperformed by PKD on other datasets. AGST is excessively dependent on the original graph topology for label propagation, rendering it vulnerable to structural noise and facing significant challenges when transferred to large-scale graphs, such as Ogbn-ArXiv. IceBerg does not perform well on heterophily graphs because its capacity to disseminate information across longer distances is hampered by the proliferation of noise edges. MTAAM shows satisfactory performance on most datasets, due to its ability to autonomously identify the most valuable knowledge from each teacher during training. FairGKD achieves runner-up results on some datasets. The poor performances of KDGA and BGNN result from their excessive sensitivity to GNN selection. MSKD is equipped with the fixed message-passing mechanism, showing that the single message-passing mechanism underperforms on all the datasets compared to PKD. The Random / # LN 5 approach refers to randomly selecting node predictions from 4 teachers, utilizing 5 labeled nodes per class to train teacher GNNs. The Voting / # LN 5 method selects the most frequently predicted label from 4 teachers as the annotation label. We can see that these two simple and intuitive strategies are defeated by PKD on all datasets.
Our PKD consistently achieves superior node classification results across all datasets, irrespective of the specific type of LLM. The few-shot node classification results after replacing Llama-3.1-8B-Instruct with Qwen2.5-7B-Instruct yang2024qwen2 and Mixtral-7B-Instruct-v0.3 jiang2024mixtral are shown in Table 2.
Table 2: Few-shot node classification accuracy (%) on eight TAGs using three different LLMs. The # LN 1, # LN 3, # LN 5 represent 1, 3, 5 labeled nodes per class, respectively. The best results are highlighted in dark gray, while the runner-up results are marked in light gray.
Methods Dataset Cornell Washington Texas Wisconsin
Amazon
Ratings
Wiki CS Pubmed Cora
[…]
Furthermore, to evaluate the quality of LLM-generated pseudo-labels, we compare the node classification performance of PKD and three baselines under different label settings (# LN 5, 48% training ratio expanded by the annotated labels and real labels, respectively). The experiments are conducted on four datasets (Cora, Wiki CS, Washington, and Wisconsin). The results are presented in Table 3. For GCNII and IceBerg, they are proposed to tackle the challenge of sparse labels, using the LLM-annotated node labels can improve their performance on all datasets. However, using the same number of real labels achieves better performance.
Table 3: Classification accuracy comparison under different label configurations. The best results are highlighted in dark gray, while the runner-up results are marked in light gray.
Models Labels configuration Cora Wiki CS Washington Wisconsin
GCNII # LN 5 77.74 56.29 64.17 60.94
48% LLM-generated labels 76.69 51.18 70.83 62.50
48% real labels 81.54 59.17 71.79 65.98
IceBerg # LN 5 76.23 84.88 67.76 41.53
48% LLM-generated labels 78.66 71.23 70.12 42.22
48% real labels 81.94 86.49 72.04 45.43
MSKD # LN 5 43.91 46.81 45.29 33.33
48% LLM-generated labels 45.89 54.06 48.17 39.50
48% real labels 51.61 62.73 50.39 41.51
PKD (ours) #\# LN 5 90.27 81.39 83.74 76.89
4.3 Ablation Study
Generally, the fine-tuned LLM using our proposed GTA prompts also demonstrates pretty zero-shot node classification performance, surpassing some semi-supervised GNNs from the values in Figure 2.
We assess the significance of the GTA prompts, DNS and $\mathcal{V}_{\mathcal{PR}}$ with the following default parameter settings: # LN = 3, $K$ = 4. Here, $K$ denotes the number of selected neighbors surrounding the node, to be annotated, within each embedding space of the teacher GNNs structure attributes. In the absence of DNS, neighbors are selected according to the adjacency matrix directly; in the non-use of $\mathcal{V}_{\mathcal{PR}}$, we expand the training data by random selection.
As shown in Table 4, the implementations of GAT prompts, DNS, and $\mathcal{V}_{\mathcal{PR}}$ result in varying degrees of performance improvement. Supported by fine-tuning with GTA prompts, the LLM’s enhanced logical reasoning ability, combined with high-quality neighboring nodes, substantially enhances zero-shot node classification capability, leading to superior classification performance improvement. Additionally, we also assess the methods without using reinforcement learning in the teacher selection process, including entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, random selection, and end-to-end learning. Their relevant results are provided in the Appendix E.2.
Table 4: Ablation study for GTA, DNS, and $\mathcal{V}_{\mathcal{PR}}$. $\Uparrow$ denotes an accuracy (%) increment. The three components play different roles in the improvement of the performance of our method.
[…]
Cora
45.02 Amazon Ratings
42.01
$\Uparrow$ 26.94
$\Uparrow$ 13.02
$\Uparrow$ 30.99
$\Uparrow$ 16.96
$\Uparrow$ 41.14
$\Uparrow$ 23.97
To assess the effectiveness of each part in the reward function (Equation (4)), we visualize the training processes of three variants in Figure 4: (a) $R_{1}$: The reward function for teacher GNN selection depends solely on the classification accuracy of the student GNN on the expanded training data; (b) $R_{2}$: In addition to classification accuracy, the reward function also incorporates the negative cross-entropy loss ($\mathcal{L}_{CE}$); (c) $R_{3}$: Building upon $R_{2}$, the reward function also includes the negative knowledge distillation loss ($\mathcal{L}_{DL}$). As shown in Figure 4, both the three parts contribute to the improved classification performance.
Figure 4: The comparison of different Rewards. When including all three parts simultaneously, our method (the curve in green) performs the best.
Figure 5: The effects of $K$ on homophily and heterophily graphs. When $K=4$, zero-shot node classification accuracy of the fine-tuned LLM is the highest on most graphs.
Table 5: The few-shot node classification accuracy (%) of PKD with different parameter-scale LLM.
Datasets Cora
Parameter scales 7B 14B 32B
PKD 90.07 90.58 91.54
Datasets Pubmed
Parameter scales 7B 14B 32B
PKD 85.96 86.64 87.16
4.4 Sensitivity Analysis
We investigate the impact of the hyper-parameter $K$ on the zero-shot node classification performance. We vary the value of $K$ within the range $\{1,2,3,4,5\}$ for homophily graphs and heterophily graphs to observe the variation in zero-shot node classification accuracy. As illustrated in Figure 5, accuracy exhibits significant fluctuations as $K$ changes. When $K=4$, the fine-tuned LLM demonstrates strong performance on most graphs.
To further explore the relationship between the parameter scale of LLM and PKD’s performance, we evaluated Qwen2.5-7B-Instruct with three different parameter scales: 7B/14B/32B parameters. The results are shown in Table 4.3. Obviously, the classification performance of PKD basically gets better with the increase of parameter scale. This is mainly related to the LLMs with larger parameter-scale have richer knowledge storage and better ability handling complex tasks.
Next, we investigate the ratios of nodes selected for annotating their labels by the LLM as a means to expand the training set. The results are given in Table 6. Increasing the expansion ratio can enhance the performance of PKD. This improvement can be attributed not only to the high-quality label annotation generated by the fine-tuned LLM, but also to the characteristic of PKD that is underpinned by selecting each node the most appropriate teacher GNN for knowledge distillation.
Furthermore, we perform the hyperparameter sensitivity analysis over the loss-weight coefficients $\alpha,\beta,\gamma,\eta$. For the sensitivity analysis of $\alpha$, we set $\beta=1$, $\gamma=1$, $\eta=0.5$. This strategy also applies to the sensitivity analysis of $\beta$ and $\gamma$. As for $\eta$, we fix $\alpha$, $\beta$, and $\gamma$ to their respective best values.The $\beta$ and $\eta$ are more sensitive than the other two hyperparameters, because $\beta$ is related to the supervision signal provided by the LLM-generated annotation labels and $\eta$ is related to the student GNN’s performance. All the results are reported in the Appendix E.3.
Table 6: The results with different node annotation ratios. The best results are highlighted in dark gray, while the runner-up results are marked in light gray.
Node annotation ratios 10% / # LN 5 20% / # LN 5 30% / # LN 5 40% / # LN 5 48% / # LN 5
[…]
4.5 Running Time
We also study the training efficiencies of PKD and all baselines. The running times on Cora are shown in Tabel 7. There is a trade-off between accuracy and time complexity. The incorporation of the LLM undoubtedly boosts the few-shot classification accuracy of GNNs on TAGs, but the training time increases. When applied to the bigger graphs, the time increase will be more obvious.
Table 7: Running time (second per epoch) of each method, including the pretraining process.
Datasets / Methods $\boldsymbol{T_{1}}$ $\boldsymbol{T_{2}}$ $\boldsymbol{T_{3}}$ $\boldsymbol{T_{4}}$ GCNII EGNN LLMGNN GAugLLM
Cora 0.006 0.197 0.247 0.035 0.014 0.366 0.630 0.402
Datasets / Methods Self-training AGST IceBerg KDGA MSKD BGNN MTAAM FairGKD PKD
Cora 0.016 0.018 0.011 3.911 2.289 1.001 3.318 4.100 7.314
5 Conclusions, Limitations & Future Work
In this work, we have proposed a preference-driven knowledge distillation (PKD) framework for few-shot node classification on TAGs, consisting of GNN-preference-driven Node Selector (GNS) and Node-preference-driven GNN Selector (NGS). Fine-tuned with our proposed GTA prompts, the refined LLM generates high-quality annotations. The GNS effectively determines nodes for the fine-tuned LLM to annotate and promotes knowledge distillation from the LLM to teacher GNNs. The NGS tailors for each node the most appropriate message-passing mechanism, promoting knowledge distillation from teacher GNNs to the student GNN. On various real-world TAGs, our method PKD outperforms almost all advanced GNNs and KD methods for few-shot node classification while using only a few node labels. One limitation of our method is that it is designed for TAGs. Moving forward, we plan to further explore more efficient mechanism of synergizing LLM and GNN to address the limitation of training efficiency as well as datasets beyond TAGs.
6 Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their valuable and constructive comments. This work was supported partially by the National Natural Science Foundation of China under Grants 62176184, 62476109, and 62206108, and the Fundamental Research Funds for the Central Universities.
Appendix A Detailed Description of Datasets
Table 8: Statistics of datasets. The Hom. ratio means 1-hop homophily ratio.
Dataset Cornell Washington Texas Wisconsin
Amazon
Ratings
Ogbn-
Arxiv
Wiki CS Pubmed Cora
Hom. ratio 0.1504 0.1545 0.1989 0.2109 0.4777 0.6542 0.6588 0.7924 0.8252
# Node 189 214 168 264 5068 169343 11701 19717 2708
# Edge 166 182 91 388 17334 1166243 216123 88648 10556
# Features 1703 1703 1703 1703 300 128 300 500 1433
# Classes 5 5 5 5 5 40 10 3 7
Domain Web page Web page Web page Web page Co-purchase Co-citation Wikipedia page Co-citation Co-citation
Cornell, Washington, Texas, Wisconsin:
These four datasets are derived from the WebKB webpage dataset, collected from the computer science departments of various universities. In these datasets, nodes represent web pages, while edges denote hyperlinks connecting them. All words from the given web pages are collected as the features for the nodes. The webpage categories can be listed as following: Student, Project, Course, Staff, Faculty.
Amazon Ratings:
This dataset is derived from the Amazon product co-purchasing network metadata, sourced from the SNAP datasets [52]. Nodes represent products (Books, Music CDs, DVDs, Videos) and edges signify relationships between products that are frequently co-purchased. The task involves predicting the average rating assigned to each product by reviewers. The possible rating values are grouped into five distinct classes. For node features, we utilize the NV-Embed-v2 [53] embeddings generated from the product descriptions. To reduce the size of the graph, we only consider the largest connected component of the 5-core of the graph.
Wiki CS:
Wiki CS is a graph derived from the Wikipedia platform. The nodes in Wiki CS represent Wikipedia page descriptions, while the edges correspond to hyperlinks between distinct pages. The Wiki CS dataset and its raw text [44] are sourced from OFA [54]. The graph consists of 11,701 nodes and 216,123 edges. The Wiki CS dataset is suitable for node classification tasks. The Wiki CS dataset is categorized into 10 distinct categories: Computational Linguistics, Databases, Operating Systems, Computer Architecture, Computer Security, Internet Protocols, Computer File Systems, Distributed Computing Architecture, Web Technology, Programming Language Topics.
Cora, Pubmed and Ogbn-Arxiv:
The Cora dataset represents a co-citation graph of computer science research papers. The dataset is sourced from OFA [54], with the original data derived from [9]. In [9], the authors recollect the dataset due to the commonly employed bag-of-words features in the widely used Cora dataset within the GNN community, where raw text is difficult to retrieve. The revised Cora dataset contains 2,708 nodes and 10,556 edges, matching the specifications of the original dataset. The dataset is divided into 7 categories: Theory, Reinforcement Learning, Genetic Algorithms, Neural Networks, Probabilistic Methods, Case-Based, Rule Learning.
The Pubmed dataset represents a co-citation graph of biomedical research papers focused on diabetes mellitus. The source and processing procedure of Pubmed are identical to those of Cora. After processing, the dataset consists of 19,717 nodes and 88,648 edges. The dataset is classified into 3 categories: Experimentally Induced Diabetes, Type 1 Diabetes, Type 2 Diabetes.
The Ogbn-Arxiv dataset is a citation graph of papers from the arXiv platform. It is collected from the Arxiv dataset and its raw text as OGB[43] and OFA [54]. There are 169,343 nodes and 1,166,243 edges in the graph. It contains 40 sub-categories of compute science.
Appendix B Detailed Prompts
We provide all specific prompt templates in the following for zero-shot node classification, Node-preference-driven GNN Selector and GTA Prompts, respectively.
B.1 Prompts for Zero-shot Node Classification
The complete prompts for zero-shot node classification are provided as below. Similarly, for each dataset, we refine specific descriptions to ensure contextual coherence.
Role Prompt
System Prompt Papers in this field can be divided into 7 categories: [Case Based, Genetic Algorithms, Neural Networks, Probabilistic Methods, Reinforcement Learning, Rule Learning, Theory]. You will serve as an assistant to help me to classify this target paper into the 7 categories above according to its description and related papers’ descriptions, who may be of the same category as this target paper. I will provide you with the descriptions of this target paper and its related papers.
Here are the instructions:
I will provide you with information in the form of a JSON string that describes the target paper:
Title: the title of this target paper. Abstract: the abstract of this target paper.
Related Title: the title of the related paper. Related Abstract: the abstract of the related paper.
Related Title: the title of the related paper. Related Abstract: the abstract of the related paper.
Requirements:
❶ Please provide your response in JSON format, following this structure:
Reasoning: Briefly explain your reasoning process for the predicted category.
Category: The best category you predict for this paper, this category must belong to these 7 categories: [Case Based, Genetic Algorithms, Neural Networks, Probabilistic Methods, Reinforcement Learning, Rule Learning, Theory];
❷ There are 2000 words limits for the reasoning;
❸ Do not provide any other text outside the JSON string;
❹ Focus only on content in the actual text and avoid making false associations;
❺ The output can only contain category and reasoning.
User Prompt Title: $\mathbf{t}_{title}$. Abstract: $\mathbf{t}_{abstract}$.
Related Title: $\mathbf{t}_{title}^{r_{1}}$. Related Abstract: $\mathbf{t}_{abstract}^{r_{1}}$.
Related Title: $\mathbf{t}_{title}^{r_{2}}$. Related Abstract: $\mathbf{t}_{abstract}^{r_{2}}$.
Related Title: $\mathbf{t}_{title}^{r_{3}}$. Related Abstract: $\mathbf{t}_{abstract}^{r_{3}}$.
B.2 Prompts for Node-preference-driven GNN Selector
Unlike the prompts used zero-shot node classification described above, we do not collect responses from the LLM; instead, we focus solely on the outputs generated by the subsequent projector. Similarly, for each dataset, we refine certain descriptions to maintain contextual consistency.
Role Prompt
System Prompt There are four names of teacher networks: [[APPNP, GCN, $\text{H}_{2}\text{GCN}$, GAT]]. We need to perform knowledge distillation for each node in this graph consist of nodes (papers) and edges (citation relationships). You will serve as an assistant to help me to assign the best teacher network for the target node (paper) based on the following information.I will provide you with three kinds of attributes of the target node (paper).
Here are the instructions:
I will provide you with information in the form of a JSON string that describes the node (paper):
Semantic attributes: the title and abstract of this paper.
Structure attributes: four teacher networks’ logit output of this target node.
Prediction attributes: important neighbors (papers), which are closely related the target node (paper) and their contents.
Requirements:
❶ Please provide your response in JSON format, following this structure:
Reasoning: Briefly explain your reasoning process for the selected teacher network.
Teacher network: The best teacher network you assign for this node (paper), this result must belong to these 4 teachers: [[APPNP, GCN, $\text{H}_{2}\text{GCN}$, GAT]];
❷ There are 2000 words limits for the reasoning;
❸ Do not provide any other text outside the JSON string;
❹ Focus only on content in the actual text and avoid making false associations;
❺ The output can only contain teacher network and reasoning.
User Prompt Semantic attributes: It is the content description of this target paper: $\mathbf{t}$.
Structure attributes: It has following important neighbors (papers), which are closely related the target paper. Their content descriptions are: …
Prediction attributes:
The APPNP’s logits output of this target paper is $\text{str}(\mathbf{z}_{APPNP})$,
The GCN’s logits output of this target paper is $\text{str}(\mathbf{z}_{GCN})$,
The $\text{H}_{2}\text{GCN}$’s logits output of this target paper is $\text{str}(\mathbf{z}_{\text{H}_{2}\text{GCN}})$,
The GAT’s logits output of this target paper is $\text{str}(\mathbf{z}_{GAT})$
B.3 Graph Topology Aware (GTA) Prompts
Generating effective prompts for graph-based tasks can be challenging for LLMs, due to the inherent complexity of graph structures and relationships that must be accurately represented. To address this challenge, we propose structured-tasks text for graph topology aware, designed specifically for fine-tuning LLMs.
Task 1: Connectivity
This task is determining whether or not two nodes $v_{i}$ and $v_{j}$ in an undirected graph are connected. Specifically, we randomly select node pairs $v_{i},v_{j}\in\mathcal{V}$ and ask whether or not an edge exists between them in the graph, answering with a ”True/False” response. To ensure prompt diversity, only one-third of the possible node pairs are selected for each graph.
Task 2: Degree
The degree of a node, $D$, is the number of nodes directly connected to it. In this task, we group nodes based on their degree and select a node $v_{i}$ from a group. The LLM is then given the node’s local structure according to the adjacency matrix $\mathbf{A}$, and is asked for the degree of the node. To prevent repetitive prompts, only one-third of the nodes from each degree group are selected.
Task 3: Cycle Detection
A cycle in an undirected graph without self-loop is a path where the first and last nodes are the same. This task requires the LLM to answer whether a cycle exists in the given sequence of nodes, $\{v_{1},...,v_{l},...,v_{1}\}$. We generate random walks [55] of length greater than 10 and arrange them into node sequences. After describing their neighbors information (derived from the adjacency matrix $\mathbf{A}$), the LLM is then asked whether or not any sequence of nodes forms a cycle.
Task 4: Text Generation
We randomly select a node set $\mathcal{W}=\{v_{i}\}_{i=1}^{N/3}$ as the source nodes, and a breadth-first search (BFS) is conducted from each source node to identify nodes in graph at a distance greater than $t$ edges from $v_{i}$, which are collected as target nodes. Redundant nodes are removed via the long-to-short path conversion module [56]. The LLM is tasked with generating textual descriptions of target nodes based on the semantic attributes of the preceding nodes in the path.
Specifically, Task 1 enhances the LLM’s ability to identify neighboring nodes and understand the structure of local neighborhoods; Task 2 strengthens the LLM’s ability to recognize the significance of node degrees within the graph context; Task 3 reinforces the LLM to reason about complex graph topologies, such as cycles and long-range node dependencies; Task 4 improves path-based reasoning and contextualization of nodes in the local graph structure.
Role Prompt
System Prompt You will serve as a graph machine learning expert in connectivity detection to help me to determine whether the edge exists between the given two targeted nodes. There is a undirected graph consisting of papers (nodes) and the citation relationships (edges) between them. I will provide the information of the two targeted nodes and their neighbors, consisting of indexes, textual content.
Here are the instructions:
I will provide you with information in the form of a JSON string that describes the target papers:
The first targeted paper:
Node index: …; Title: …; Abstract: …;
The $k_{th}$ neighbor: Index:…; Title: …; Abstract: …;
The second targeted paper:
Node index: …; Title: …; Abstract: …;
The $k_{th}$ neighbor: Index:…; Title: …; Abstract: …;
Requirements:
❶ Please provide your response in JSON format, following this structure:
Reasoning: Briefly explain your reasoning process for the selected teacher network.
Answer: You only can select one from [[True, False]] as the best answer;
❷ There are 2000 words limits for the reasoning;
❸ Do not provide any other text outside the JSON string;
❹ Focus only on content in the actual text and avoid making false associations;
❺ The output can only contain answer and reasoning.
User Prompt The first targeted paper:
Node index: $i$; Title: $\mathbf{t}_{title}^{i}$; Abstract: $\mathbf{t}_{abstract}^{i}$
The $k_{th}$ neighbor’s node index: $I_{k}^{i}$ Title: $\mathbf{t}_{title}^{I_{k}^{i}}$ Abstract: $\mathbf{t}_{abstract}^{I_{k}^{i}}$
The second targeted paper:
Node index: $j$; Title: $\mathbf{t}_{title}^{j}$; Abstract: $\mathbf{t}_{abstract}^{j}$
The $k_{th}$ neighbor’s node index: $I_{k}^{j}$ Title: $\mathbf{t}_{title}^{I_{k}^{j}}$ Abstract: $\mathbf{t}_{abstract}^{I_{k}^{j}}$
The full prompts for Connectivity is presented below. When generating prompts for different datasets, we adjust certain descriptions to better align with the specific context. For example, when constructing prompts for Texas, the background description should be adapted to reflect web pages, and the relationship should be revised to hyperlinks, along with other context-specific adjustments. Similarly, for each task, the prompts must also be modified to correspond to the specific content described in Section 3.2.
Appendix C Implementation Details and Time Complexity Analysis
Input: $\mathcal{G}_{T}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{A},\mathbf{T})$, training dataset with true labels $\mathcal{D}_{L}$, teacher GNNs $\{T_{b}\}_{b=1}^{4}$ with parameters $\{f_{T_{b}}^{\theta}\}_{b=1}^{4}$, student GNN $S$ with parameter $f_{S}^{\theta}$, fine-tuned LLM $LLM^{\theta}$, Policy Model $f_{A}^{\theta}$, Value Model $f_{V}^{\phi}$, epoch number of RL $L_{1}$
Output: The expanded training dataset $\tilde{\mathcal{D}_{L}}$, optimized parameters $LLM^{\theta^{\ast}},f_{S}^{\theta^{\ast}},f_{A}^{\theta^{\ast}}$, $f_{V}^{\phi^{\ast}}$ and predicted labels $\tilde{y}$.
1 $\tilde{\mathcal{D}_{L}}\leftarrow\mathcal{D}_{L}$ // Training dataset initialization ;
// LLM fine-tuning;
2 $LLM^{\theta^{\ast}}\leftarrow LLM^{\theta}$;
// GNN-preference-driven Node Selector;
3 Filter out GNN-preference nodes based on the preference rank $\mathcal{V}_{\mathcal{PR}}$ and get their annotations from $LLM^{\theta^{\ast}}$;
4 Conduct prediction distillation from $LLM^{\theta^{\ast}}$ to $\{f_{T_{b}}^{\theta}\}_{b=1}^{4}$ for retrain them ;
// Node-preference-driven GNN Selector;
5 for $l_{1}\leftarrow 1$ to $L_{1}$ do
6   Shuffle $\tilde{\mathcal{D}_{L}}$ to get a new training sequence;
7   Complete prompts $\{\mathcal{P}_{i}\}_{i=1}^{W}$ for each selected nodes;
8 for each node $v_{\mathcal{PR}}\in\tilde{\mathcal{D}_{L}}$ do
9      NSG select teacher GNN for $v_{\mathcal{PR}}$ and get one-hot vector $\mathbf{m}_{i}$;
10      Update the parameter $f_{S}^{\theta}$ and get reward $R_{i}$ by Eqn. (1);
11      Store ($\mathcal{P}_{i},\mathbf{m}_{i},R_{i}$) to the episode history $\mathcal{F}$;
12
13  Update the parameter $f_{A}^{\theta}$ and $f_{V}^{\phi}$ by Eqn. (5) and Eqn. (6) ;
14
15return $\tilde{\mathcal{D}_{L}},LLM^{\theta^{\ast}},f_{S}^{\theta^{\ast}},f_{A}^{\theta^{\ast}},f_{V}^{\phi^{\ast}}$;
Algorithm 1 The training of PKD.
First of all, we outline the training setup employed for the experiments detailed in Section 4.2. Uniform training hyper-parameters are applied across all baseline models and datasets. Specifically, the following hyper-parameter values are utilized: the hidden dimension is set to 128. We use ReLU activation functions in all our baseline models. The Adam optimizer is utilized with a learning rate of $1\times e^{-2}$ and weight decay of $5\times e^{-4}$. We train each baseline for 600 steps and select the best step based on the validation accuracy. In our proposed method, we train the student 5 epochs after GNN selection driven by node attributes every time and train the agent 200 epochs. The other weight hyper-parameters are set as follows: $\alpha=0.5,\beta=1,\gamma=0.1,\eta=0.3,c_{1}=0.5,c_{2}=0.01,\epsilon=0.2$.
Additionally, the parameters of Action Model and Value Model are updated as follows:
$f_{A}^{\theta}\leftarrow f_{A}^{\theta}-\rho_{A}\nabla_{f_{A}^{\theta}}(\mathcal{L}_{A}+c_{1}\mathcal{L}_{V}-c_{2}H(\bm{\pi}_{T}))$ (5)
$f_{V}^{\phi}\leftarrow f_{V}^{\phi}-\rho_{V}\nabla_{f_{V}^{\phi}}\mathcal{L}_{V}$ (6)
where $f_{A}^{\theta}$ and $f_{V}^{\phi}$ represent the trainable parameters of the Policy Model and Value Model, respectively. $\rho_{A}$ and $\rho_{V}$ are their learning rates and $\nabla_{f_{A}^{\theta}}$ and $\nabla_{f_{V}^{\phi}}$ are the gradients of their parameters. $\mathcal{L}_{A}$ and $\mathcal{L}_{V}$ are objective functions belonging to the Policy Model and Value Model, respectively. $c_{1},c_{2}$ are hyper-parameters to balance weights. $H(\bm{\pi}_{T})$ is employed to enhance the entropy of the policy and promote sufficient exploration. Based on the CLIP strategy [41], the final objective function of the Policy Model is:
$\mathcal{L}_{A}=-\mathbb{E}_{i}[\text{min}(r_{i}(f_{A}^{\theta})\hat{A}_{i},\text{clip}(r_{i}(f_{A}^{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{i})]$ (7)
where $\mathbb{E}_{i}$ represents the expectation in the time step $i$. $r_{i}(f_{A}^{\theta})$ is the ratio of the $i$-th policy to the (i1)(i-1)-th policy. $\hat{A}_{i}$ is the advantage estimation in the current step, denoting how good or bad the Action is. $\epsilon$ is a hyper-parameter, which determines the range of the CLIP operation.
The objective functions of the Value Model and $H(\bm{\pi}_{T})$ are:
$\mathcal{L}_{V}=\mathbb{E}_{i}[(f_{V}^{\phi}(\mathcal{P}_{i})-\hat{R}_{i})^{2}]$ (8)
$H[\bm{\pi}_{T}]=-\mathbb{E}_{i}[\pi_{f_{A}^{\theta}}(A_{T}|\mathcal{P}_{i})\text{log}\pi_{f_{A}^{\theta}}(A_{T}|\mathcal{P}_{i})]$ (9)
where $f_{V}^{\phi}(\mathcal{P}_{i})$ and $\hat{R}_{i}$ denote the Value Model’s estimation of State $\mathcal{P}_{i}$ and the target value of real Reward $R_{i}$, respectively. $A_{T}$ denotes the specific action and $\pi_{f_{A}^{\theta}}(A_{T}|\mathcal{P}_{i})$ is the probability that Policy $f_{A}^{\theta}$ takes action $A_{T}$ in state $\mathcal{P}_{i}$.
The detailed training procedure is shown in Algorithm 1.
The specific analysis of the time complexity of PKD training and testing are provided below:
The time complexity of PKD training is mainly divided into three parts: LLM fine-tuning, GNN-preference-driven Node Selector and Node-preference-driven GNN Selector. The GNN-preference-driven Node Selector also can be divided into the annotations generation and prediction distillation.
Table 12: The GTA fine-tuning configurations on Llama-3.1-8B-Instruct.
Dataset size Epoch lora_r lora_alpha optimizer learning rate GTA fine-tuning time LLM
53,617 2 4 4 AdamW $1e^{-4}$ 9h 41m 48s Llama-3.1-8B-Instruct
First, we use Low-Rank Adaptation (LoRA) strategy [57] for efficient parameter training, with hyperparameters set to $r=4$, $\alpha=4$, $epoch=2$ (as shown in Table 12), and the rest are set according to the default settings of llama-factory [^1] https://github.com/hiyouga/LLaMA-Factory. Weight merge is also involved. In general, the time complexity of this part is $\mathcal{O}(nLdr+Ld^{2}r)$, where $n$ is the number of instructions, $L$ is the number of layers applying Lora, and d is the dimension of the LLM hidden layer. $r\ll d$, so the time complexity is bound by $\mathcal{O}(NLd+Ld^{2})$.
The process of annotations generation includes sorting the selected nodes and the reasoning process of LLM, and its time complexity is $\mathcal{O}(W\log W)$ and $\mathcal{O}(WL^{\prime}(l^{2}d+ld^{2}))$, where $W$ is the number of selected nodes, $L^{\prime}$ is the number of transformer layers in LLM, and $l$ is the input sequence length. Generally, $W\ll l$, $L^{\prime}\ll l$, then the time complexity is $\mathcal{O}(l^{2}d+ld^{2})$.
The time complexity of teacher GNN (2-layers) re-training is bound by $\mathcal{O}((NF+M)D)$, where $N$ is the number of nodes, $F$ is the node feature dimension, $M$ is the number of edges, and $D$ is the GNN hidden layer dimension.
The time complexity of Node-preference-driven GNN Selector is $\mathcal{O}(W(l^{2}d+ld^{2}+dd’+d^{\prime}a))$, where $d^{\prime}$ is the dimension of the MLP hidden layer, $a$ is the number of action categories, $W\ll l$, $a\ll d$, so the time complexity is bound by $\mathcal{O}(l^{2}d+ld^{2}+dd^{\prime})$. The training time complexity of student GNN is $\mathcal{O}((NF+M)D)$ Therefore, the overall time complexity is bound by $\mathcal{O}((L+2l)d^{2}+(d’+nL)d+2l^{2}+2D(NF+M))$.
The inference time complexity of PKD is determined by the testing process of the student GNN. So its time complexity is bound by $\mathcal{O}((NF+M)D)$.
Specifically, We implement our proposed PKD with PyTorch (2.5.1) [58], PyTorch Geometric (2.6.1) [59], Python (3.10.16), Transformers (4.50.3), and vllm (0.7.0). We conduct all experiments on the NVIDIA A800-SXM4-80GB GPU and Intel(R) Xeon(R) CPU Max 9468.
The time costs for GTA fine-tuning, the Distance-based Neighbor Selector, LLM annotation, teacher re-training, and the PPO loop on Cora and Ogbn-Arixv are presented in Tables 12, 13, 14 and 15, respectively. The peak memories of performing PKD on the Cora and Ogbn-Arxiv are also listed in Table 15.
Table 13: The time costs on Cora and Ogbn-Arxiv of different LLMs.
Dataset Llama-3.1 Qwen-2.5
Parameters 8B 13B 7B 14B
Cora 0.11 GPU-hours 0.36 GPU-hours 0.05 GPU-hours 0.29 GPU-hours
Ogbn-Arxiv 1.68 GPU-hours 5.01 GPU-hours 0.94 GPU-hours 2.95 GPU-hours
Table 14: The time costs of retraining teacher GNNs on Cora, Amazon Ratings, and Ogbn-Arxiv.
Datasets Teacher GNNs Epochs Total running time (seconds)
Cora GCN, GAT, APPNP, H2GCN 300 2.4716
Obgn-Arxiv GCN, GAT, APPNP, H2GCN 300 18.2017
Amazon Ratings DirGNN, GPRGNN, HoloNets, H2GCN 300 3.1452
Table 15: Running times of PPO on Cora and Obgn-Arixv. "m" and "s" denote minute and second.
Dataset Peak memory Running time / epoch
Cora 454.62 MB 7.9s
Obgn-Arixv 1655.18 MB 44m 8.3s
Appendix D Proofs for Propositions 3.1
The uncertainty usually refers to a measure of the confidence of a model in predicting a certain sample. From the perspective of collective consensus [40], we define the $K$-uncertainty of one node as the deviation of each teacher GNN’s prediction probability distribution from the overall prediction probability distribution. From the Proposition 3.1, we can get that, the larger $\delta_{K}$ of one node, the stronger the uncertainty of this node, which is more beneficial to teacher GNNs training.
Proof.
For each node $v$, the prediction probability distributions of $B$ teacher GNNs can be denoted by $P_{1},P_{2},...,P_{B}$. The $K$-uncertainty of node $v$ is defined as:
$\delta_{K}(v)\triangleq\sum_{1\leq i<j\leq B}^{N}[D_{KL}(P_{i}(v)||P_{j}(v))+D_{KL}(P_{j}(v)||P_{i}(v))]$ (10)
Here, we define the average prediction probability distribution as following:
Definition D.1.
The average prediction probability distribution $\mathcal{M}$ is the benchmark for the overall prediction probability distribution to measure both the models confidence and the consistency of each GNN with the overall probability distribution.
$\mathcal{M}(v)=\frac{1}{B}\sum_{i=1}^{B}P_{i}(v)$ (11)
Then, the uncertainty of node $v$ is,
$\delta_{v}=\frac{1}{B}\sum_{i=1}^{B}D_{KL}(P_{i}(v)||\mathcal{M}(v))$ (12)
According to Jenson’s inequality, we have
$\delta_{K}(v)\geq N\delta_{v}$ (13)
For any probability distribution $P$, we have
$D_{KL}(P||\mathcal{M})=H(P,\mathcal{M})-H(P)$ (14)
Then,
$\delta_{K}=\frac{1}{N}\sum[H(P_{i},\mathcal{M})-H(P_{i})]$ (15)
As the $K$-uncertainty increases, the entropy of the GNN’s prediction probability distribution $P_{i}$ increases, and the cross-entropy $H(P_{i},\mathcal{M})$ grows significantly due to the larger probability distribution differences. So there is,
$\delta_{K}(v)\varpropto\sum_{i=1}^{B}[H(P_{i},\mathcal{M})-H(P_{i})]\varpropto\delta_{v}$ (16)
That is,
$\delta_{K}(v)\varpropto\delta_{v}$ (17)
From the Proposition 3.1, we also can get that, selecting high-uncertainty nodes to expand the training set benefits GNNs training.
For a GNN with prediction probabilities $P(y=c|v)$, the entropy of an unlabeled node $v$ is
$H(v)=-\sum_{c=1}^{C}P(y=c|v)\log P(y=c|v)$ (18)
To maximize information gain, we select the node $v^{*}$ with the highest uncertainty (entropy):
$v^{*}=\mathop{\arg\max}\limits_{v}H(v)$ (19)
After expanding $v^{*}$ to the training dataset, the loss function becomes:
$\mathcal{L}_{new}=\mathcal{L}_{old}(\theta)+\mathcal{L}(f_{\theta}(v^{*}),y^{*})$ (20)
Here, $y^{*}$ is considered the true label based on the fine-tuned LLM. The GNN parameters are updated as:
$\theta_{new}=\theta_{old}-r\cdot\nabla_{\theta}\mathcal{L}(f_{\theta}(v^{*}),y^{*})$ (21)
Since the prediction probability distribution of $v^{*}$ is close to uniform (due to high entropy) [60], the gradient more effectively corrects the GNN parameters [61], reducing the error. According to the preference rank: $\mathcal{V}_{\mathcal{PR}}=\text{Sort}({\left\{v_{1},\ldots,v_{N}\right\},\delta_{K}(v_{1}),\delta_{K}(v_{2}),\ldots,\delta_{K}(v_{N})})$, we can get the follows:
${f_{T}^{\theta}}^{*}(\tilde{\mathcal{D}}_{L})=\mathop{\arg\min}\limits_{v_{i}\in\{v_{\mathcal{PR}}^{1},v_{\mathcal{PR}}^{2},\ldots,v_{\mathcal{PR}}^{W}|\delta_{K}(v_{\mathcal{PR}}^{W})>\tilde{\delta}_{K}\}}\frac{1}{W}\sum\mathcal{L}(f_{T}^{\theta},v_{i})$ (22)
where $\tilde{\mathcal{D}}_{L}$ is the expanded training dataset. $f_{T}^{\theta}$$^{*}$ is the optimal parameter of teacher GNN. $v_{\mathcal{PR}}^{w}$ represents the $w$-th nodes in the preference rank.
Appendix E Other Experimental Results
E.1 Visualization
Figure 6 presents the outstanding node classification performance we mentioned in Section 4.2, which is illustrated by the t-SNE [62] visualization of the embedding spaces for Cora. Notably, Figure 6(a) illustrates the results of the student GNN (GCN) under the # LN 5 condition.
From the Figure 6, we can see that, some KD methods fail to enable the student GNN to learn discriminative node representations, as evidenced by the absence of clustered structures in the embedding space, exemplified by MSKD, BGNN, and MTAAM. GCNII and KDGA struggle to form well-separated clusters, whereas methods like LLMGNN, GAugLLM, and FairGKD yield clusters with limited purity. Compared to these baselines, our method generates embeddings with significantly enhanced inter-class separability and high cluster purity, resulting in improved few-shot node classification performance.
(a) Original student GNN
(b) GCNII
(c) EGNN
(d) LLMGNN
(e) GAugLLM
(f) AGST
(g) KDGA
(h) MSKD
(i) BGNN
(j) MTAAM
(k) FairGKD
(l) PKD # LN 5
Figure 6: t-SNE [62] visualizations on Cora.
E.2 Ablation Study
The results on Cora and Amazon Ratings of our RL-based method (PKD) and the other three teacher selection methods (entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, random selection, and end-to-end learning) are shown in Table 16. It is obvious that the RL-based method significantly outperforms the other three methods.
Table 16: A comparison of Entropy-based ranking, Random selection, End-to-end learning and RL-based approach on Cora and Amazon Ratings.
Methods Cora Amazon Ratings
Entropy-based ranking 75.70 55.74
Random selection 62.80 63.05
End-to-end learning 60.29 60.39
RL-based (ours) 90.27 65.93
E.3 Hyperparameters Sensitivity Analysis
As mentioned in Sec. 4.4, we perform the hyperparameter sensitivity analysis over the loss-weight coefficients $\alpha,\beta,\gamma,\eta$ on two datasets, and report the results in Tables 17 18 19 20. As result, the proposed PKD can achieve much better performance when $\alpha=0.5,\beta=1,\gamma=0.1,\eta=0.3$.
Table 17: The influence of $\alpha$.
$\alpha$ 2 1 0.9 0.8 0.7 0.6 0.5 0.4 0.3
Cora 70.12 69.49 72.71 69.05 72.41 73.15 75.66 74.70 73.78
Amazon Ratings 61.74 61.66 60.47 60.28 61.30 62.76 63.83 62.86 61.81
Table 18: The influence of $\beta$.
$\beta$ 2 1 0.5 0.25
Cora 74.15 75.36 67.17 64.10
Amazon Ratings 63.93 63.91 63.23 62.94
Table 19: The influence of $\gamma$.
$\gamma$ 2 1 0.5 0.2 0.1 0.05
Cora 62.66 65.06 64.40 63.77 70.12 68.90
Amazon Ratings 63.93 63.93 63.63 63.71 64.01 62.58
Table 20: The influence of $\eta$.
$\eta$ 0.9 0.7 0.5 0.3 0.1
Cora 89.42 86.67 88.43 90.27 89.96
Amazon Ratings 63.81 63.75 64.79 65.93 64.87
Appendix F Broader Impact
The proposed PKD offers significant broader impacts by enhancing few-shot node classification on TAGs. By combining the strengths of LLM and GNN, it improves learning efficiency, reducing the need for expensive and time-consuming manual annotation. This can benefit industries like social media, recommendation systems, and network analysis, enabling more accurate and scalable models for personalized services, fraud detection, and dynamic optimization.
Additionally, PKD can tailor message-passing mechanisms to node-specific attributes can lead to more adaptive and efficient machine learning models. It also democratizes access to advanced machine learning, allowing smaller organizations and researchers with limited resources to develop effective models. However, ethical considerations, such as privacy and fairness, must be prioritized to ensure responsible deployment.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In this paper, we propose a preference-driven knowledge distillation (PKD) framework that unites LLMs and various-architectures GNNs for few-shot node classification on TAGs. We claim the contributions and scope in the abstract and introduction sections (See Abstract and Introduction Section).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In this work, we discuss the limitations of our research and outline directions for future work (See Conclusion).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In this work, we provide the Proposition 3.1 and its complete proof(See Method and Appendix. D).
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the code necessary for replicating the studies described in this paper via an anonymous link, and we detail the experimental setup for the replication in the article itself (See Experiments and Appendix. C).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: For the datasets disclosed in the article, we have provided information regarding their sources and origins (See Appendix. A).
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: we have specified all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results (See Experiments and Appendix. C).
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In this paper, we have reported the standard deviation of the experiments (See Experiments and Appendix. E).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In this paper, we provide detailed information about the experimental resources, including GPU configurations used in our studies and running time costs about all methods (See Experiments).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The study presented in this paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have provided the societal impacts of the work (See Appendix F).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not address issues related to this aspect.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All creators and original owners of the assets used in our paper, such as code, data, and models, have been properly credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The research presented in this paper is not concerned with new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve experiments or research related to human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not address potential risks incurred by study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLMs is an important component of the core methods in this research and we has describe the usage in detail (See Method, Experiments and Appendix C).
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
 Graph neural networks (GNNs) kipf2016semi ; yun2019graph have demonstrated excellent performance and efficiency in node classification on TAGs, which are supported by high-quality labels and effective message-passing mechanisms zhu2021textgnn .
-However, the manual labeling of nodes is undoubtedly a tedious, expensive, and time-consuming task shi2025latent ; cao2025low .
+However, the manual labeling of nodes is undoubtedly a tedious, expensive, and time-consuming task shi2025latent ; cao2025analytical .
 In many scenarios, only a few node labels are available.
 Additionally, nodes often have complex and diverse interaction relationships with each other—their local topologies are intricate—which challenge traditional GNNs with fixed message-passing mechanisms.
-Compared with GNNs, large language models (LLMs) exhibit impressive zero-/few-shot learning capabilities on TAGs chen2024exploring ; wang2024llms ; wu2024exploring .
-But the large parameter scale considerably hinders their inference efficiency he2023harnessing .
+Compared with GNNs, large language models (LLMs) exhibit impressive zero-/few-shot learning capabilities on TAGs chen2024exploring ; wang2024llms ; wu2025exploring .
+But the large parameter scale considerably hinders their inference efficiency he2024harnessing .
 A natural idea is to blend their complementary strengths for few-shot node classification on TAGs.
@@
 However, directly distilling knowledge from the LLM to GNN is impractical.
-Firstly, the discrepancy of decoder-only (LLMs) and encoder-only (GNNs) leads to fundamentally different characteristics in their embedding spaces chen2023label .
+Firstly, the discrepancy of decoder-only (LLMs) and encoder-only (GNNs) leads to fundamentally different characteristics in their embedding spaces chenlabel .
 And the huge embedding-dimension difference needs sophisticated embedding alignment and also brings high training cost hu2025large .
@@
 It regards the fine-tuned LLM as the RL-based (reinforcement learning) agent, which treats all textualized node-specific attributes (including node’s semantic, structure, and prediction attributes) as state and the student GNN’s performance as reward.
-Contributions
-❶We introduce a preference-driven knowledge distillation (PKD) framework to synergize the complementary strengths of the LLM and various GNNs ingeniously for few-shot node classification on TAGs; ❷ we propose a GNN-preference-driven node selector, effectively determining nodes for annotation by the LLM and promoting knowledge distillation from the LLM to teacher GNNs; ❸ we propose a node-preference-driven GNN selector to tailor for each node the most appropriate message-passing mechanism, promoting knowledge distillation from teacher GNNs to the student GNN; ❹ we validate the efficacy of PKD for few-shot node classification on nine TAGs.
-The experiments show that it even defeats some state-of-the-art methods that use more node labels.
+Our contributions can be summarized as follows:
+  • We introduce a preference-driven knowledge distillation (PKD) framework to synergize the complementary strengths of the LLM and various GNNs ingeniously for few-shot node classification on TAGs;
+  • We propose a GNN-preference-driven node selector, effectively determining nodes for annotation by the LLM and promoting knowledge distillation from the LLM to teacher GNNs;
+  • We propose a node-preference-driven GNN selector to tailor for each node the most appropriate message-passing mechanism, promoting knowledge distillation from teacher GNNs to the student GNN;
+  • We validate the efficacy of PKD for few-shot node classification on nine TAGs. The experiments show that it even defeats some state-of-the-art methods that use more node labels.
 2 Related Work
 2.1 Graph Neural Networks
-The field of graph learning has been dominated by GNNs, such as GCN kipf2016semi , GAT velivckovic2017graph , APPNP gasteiger2018predict , $\text{H}_{2}\text{GCN}$ zhu2020beyond , and GPRGNN chien2020adaptive .
-Recently, HoloNets koke2023holonets introduces a dual-filter mechanism with spectral response, extending spectral convolutions to directed graphs.
+The field of graph learning has been dominated by GNNs.
+Early GCN kipf2016semi introduces a spectral-based graph convolution operation to propagate node information through the graph.
+GAT velivckovic2018graph uses attention mechanisms to weigh neighbors’ contributions, enabling adaptive learning of neighborhood importance.
+APPNP gasteiger2018combining enhances message passing by using personalized propagation with a power iteration approach, improving label propagation on graphs.
+$\text{H}_{2}\text{GCN}$ zhu2020beyond extends GCNs by incorporating higher-order neighborhood information to improve representation power.
+GPRGNN chien2021adaptive combines graph convolution with residual connections to improve propagation efficiency, particularly in graphs with diverse node degrees.
+HoloNets koke2024holonets introduces a dual-filter mechanism with spectral response, extending spectral convolutions to directed graphs.
 DirGNN rossi2024edge defines the in-neighbors and out-neighbors and performs separate propagation and aggregation, improving the message passing through the incorporation of edge directionality.
@@
 DMKD liu5084903distilling harnesses complementary knowledge from various GNNs and conducts layer-level knowledge distillation to mitigate the constraint of a single teacher.
-Furthermore, chen2023label is a label-free method that proposes the LLM-GNN.
+Furthermore, chenlabel is a label-free method that proposes the LLM-GNN.
 It uses LLMs to get high-quality annotation through active and confidence-awareness node selection, thereby circumventing the difficulty of label annotation by humans.
@@
 Datasets
-In order to assess the few-shot node classification performance of our method on TAGs, we conduct a comprehensive series of experiments across 9 real-world datasets: Cornell, Washington, Texas, Wisconsin zhu2020beyond , Amazon Ratings platonov2023critical , Ogbn-Arxiv hu2020open , Wiki CS mernyei2020wiki , Pubmed, Cora yang2016revisiting .
-They have various 1-hop homophily ratios yang2022incorporating and additional details of the datasets can be found in Appendix A.
-For the KD-baselines, we partition the nodes of each graph into training, validation, and test sets, allocating 48%, 32%, and 20%, respectively, based on the proportion division mentioned in liang2024sign .
+In order to assess the few-shot node classification performance of our method on TAGs, we conduct a comprehensive series of experiments across 9 real-world datasets: Cornell, Washington, Texas, Wisconsin zhu2020beyond , Amazon Ratings platonov2023a , Ogbn-Arxiv hu2020open , Wiki CS mernyei2020wiki , Pubmed, Cora yang2016revisiting .
+They have various 1-hop homophily ratios yang2024incorporating and additional details of the datasets can be found in Appendix A.
+For the KD-baselines, we partition the nodes of each graph into training, validation, and test sets, allocating 48%, 32%, and 20%, respectively, based on the proportion division mentioned in pmlr-v235-liang24c .
 For PKD and other baselines, we randomly select 1, 3, and 5 labeled nodes per class as the initial training data and then expand the dataset to 48% of the total using the GNS module.
@@
 Baselines
-We compare our method against the following baseline models: (i) Advanced GNNs: GCNII chen2020simple and EGNN satorras2021n ; (ii) GNNs enhanced by LLMs: LLMGNN chen2023label and GAugLLM fang2024gaugllm ; (iii) self-training for graph learning: Self-training li2018deeper , AGST ding2024toward and IceBerg li2025iceberg ; (iv) Knowledge Distillation (KD) for GNNs: KDGA wu2022knowledge , MSKD zhang2022multi , BGNN guo2023boosting , MTAAM yang2024two , and FairGKD zhu2024devil .
-For homophily graphs, the teacher GNNs used are: GCN kipf2016semi ($T_{1}$), GAT velivckovic2017graph ($T_{2}$), APPNP gasteiger2018predict ($T_{3}$), $\text{H}_{2}\text{GCN}$ zhu2020beyond ($T_{4}$), and the student is GCN; for heterophily graphs, the teacher GNNs employed are: DirGNN rossi2024edge ($T_{1}$), GPRGNN chien2020adaptive ($T_{2}$), HoloNets koke2023holonets ($T_{3}$), $\text{H}_{2}\text{GCN}$ ($T_{4}$) and the student is $\text{H}_{2}\text{GCN}$.
+We compare our method against the following baseline models: (i) Advanced GNNs: GCNII chen2020simple and EGNN satorras2021n ; (ii) GNNs enhanced by LLMs: LLMGNN chenlabel and GAugLLM fang2024gaugllm ; (iii) self-training for graph learning: Self-training li2018deeper , AGST ding2024toward and IceBerg li2025iceberg ; (iv) Knowledge Distillation (KD) for GNNs: KDGA wu2022knowledge , MSKD zhang2022multi , BGNN guo2023boosting , MTAAM yang2024two , and FairGKD zhu2024devil .
+For homophily graphs, the teacher GNNs used are: GCN kipf2016semi ($T_{1}$), GAT velivckovic2018graph ($T_{2}$), APPNP gasteiger2018combining ($T_{3}$), $\text{H}_{2}\text{GCN}$ zhu2020beyond ($T_{4}$), and the student is GCN; for heterophily graphs, the teacher GNNs employed are: DirGNN rossi2024edge ($T_{1}$), GPRGNN chien2021adaptive ($T_{2}$), HoloNets koke2024holonets ($T_{3}$), $\text{H}_{2}\text{GCN}$ ($T_{4}$) and the student is $\text{H}_{2}\text{GCN}$.
 The LLM used in the experiments is Llama-3.1-8B-Instruct dubey2024llama .
@@
 Ratings
+Ogbn-
+Arxiv
 Wiki CS Pubmed Cora
@@
 48% real labels 51.61 62.73 50.39 41.51
-PKD (ours) #\# LN 5 90.27 81.39 83.74 76.89
+PKDLlama # LN 5 90.27 81.39 83.74 76.89
 4.3 Ablation Study
@@
 In the absence of DNS, neighbors are selected according to the adjacency matrix directly; in the non-use of $\mathcal{V}_{\mathcal{PR}}$, we expand the training data by random selection.
-As shown in Table 4, the implementations of GAT prompts, DNS, and $\mathcal{V}_{\mathcal{PR}}$ result in varying degrees of performance improvement.
-Supported by fine-tuning with GTA prompts, the LLM’s enhanced logical reasoning ability, combined with high-quality neighboring nodes, substantially enhances zero-shot node classification capability, leading to superior classification performance improvement.
-Additionally, we also assess the methods without using reinforcement learning in the teacher selection process, including entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, random selection, and end-to-end learning.
-Their relevant results are provided in the Appendix E.2.
 Table 4: Ablation study for GTA, DNS, and $\mathcal{V}_{\mathcal{PR}}$.
@@
 $\Uparrow$ 23.97
-To assess the effectiveness of each part in the reward function (Equation (4)), we visualize the training processes of three variants in Figure 4: (a) $R_{1}$: The reward function for teacher GNN selection depends solely on the classification accuracy of the student GNN on the expanded training data; (b) $R_{2}$: In addition to classification accuracy, the reward function also incorporates the negative cross-entropy loss ($\mathcal{L}_{CE}$); (c) $R_{3}$: Building upon $R_{2}$, the reward function also includes the negative knowledge distillation loss ($\mathcal{L}_{DL}$).
+Figure 4: The comparison of different Rewards. When including all three parts simultaneously, our method (the curve in green) performs the best.
+As shown in Table 4, the implementations of GAT prompts, DNS, and $\mathcal{V}_{\mathcal{PR}}$ result in varying degrees of performance improvement.
+Supported by fine-tuning with GTA prompts, the LLM’s enhanced logical reasoning ability, combined with high-quality neighboring nodes, substantially enhances zero-shot node classification capability, leading to superior classification performance improvement.
+Additionally, we also assess the methods without using reinforcement learning in the teacher selection process, including entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, random selection, and end-to-end learning.
+Their relevant results are provided in the Appendix E.2.
+To assess the effectiveness of each part in the reward function (Eqn.
+(4)), we visualize the training processes of three variants in Figure 4: (a) $R_{1}$: The reward function for teacher GNN selection depends solely on the classification accuracy of the student GNN on the expanded training data; (b) $R_{2}$: In addition to classification accuracy, the reward function also incorporates the negative cross-entropy loss ($\mathcal{L}_{CE}$); (c) $R_{3}$: Building upon $R_{2}$, the reward function also includes the negative knowledge distillation loss ($\mathcal{L}_{DL}$).
 As shown in Figure 4, both the three parts contribute to the improved classification performance.
-Figure 4: The comparison of different Rewards. When including all three parts simultaneously, our method (the curve in green) performs the best.
 Figure 5: The effects of $K$ on homophily and heterophily graphs. When $K=4$, zero-shot node classification accuracy of the fine-tuned LLM is the highest on most graphs.
-Table 5: The few-shot node classification accuracy (%) of PKD with different parameter-scale LLM.
-Datasets Cora
-Parameter scales 7B 14B 32B
-PKD 90.07 90.58 91.54
-Datasets Pubmed
-Parameter scales 7B 14B 32B
-PKD 85.96 86.64 87.16
 4.4 Sensitivity Analysis
@@
 This is mainly related to the LLMs with larger parameter-scale have richer knowledge storage and better ability handling complex tasks.
+Table 5: The few-shot node classification accuracy (%) of PKD with different parameter-scale LLM.
+Datasets Cora
+Parameter scales 7B 14B 32B
+PKD 90.07 90.58 91.54
+Datasets Pubmed
+Parameter scales 7B 14B 32B
+PKD 85.96 86.64 87.16
 Next, we investigate the ratios of nodes selected for annotating their labels by the LLM as a means to expand the training set.
@@
 Cora 0.006 0.197 0.247 0.035 0.014 0.366 0.630 0.402
-Datasets / Methods Self-training AGST IceBerg KDGA MSKD BGNN MTAAM FairGKD PKD
+Datasets / Methods Self-training AGST IceBerg KDGA MSKD BGNN MTAAM FairGKD PKDLlama
 Cora 0.016 0.018 0.011 3.911 2.289 1.001 3.318 4.100 7.314
@@
 Domain Web page Web page Web page Web page Co-purchase Co-citation Wikipedia page Co-citation Co-citation
-Cornell, Washington, Texas, Wisconsin:
+Cornell, Washington, Texas, and Wisconsin:
 These four datasets are derived from the WebKB webpage dataset, collected from the computer science departments of various universities.
@@
 The Wiki CS dataset is categorized into 10 distinct categories: Computational Linguistics, Databases, Operating Systems, Computer Architecture, Computer Security, Internet Protocols, Computer File Systems, Distributed Computing Architecture, Web Technology, Programming Language Topics.
-Cora, Pubmed and Ogbn-Arxiv:
+Cora, Pubmed, and Ogbn-Arxiv:
 The Cora dataset represents a co-citation graph of computer science research papers.
@@
 The complete prompts for zero-shot node classification are provided as below. Similarly, for each dataset, we refine specific descriptions to ensure contextual coherence.
+Table 9: The prompt template for zero-shot node classification.
 Role Prompt
@@
 Similarly, for each dataset, we refine certain descriptions to maintain contextual consistency.
+Table 10: The prompt template for Node-preference-driven GNN Selector.
 Role Prompt
@@
 Specifically, Task 1 enhances the LLM’s ability to identify neighboring nodes and understand the structure of local neighborhoods; Task 2 strengthens the LLM’s ability to recognize the significance of node degrees within the graph context; Task 3 reinforces the LLM to reason about complex graph topologies, such as cycles and long-range node dependencies; Task 4 improves path-based reasoning and contextualization of nodes in the local graph structure.
+Table 11: The prompt template for the Task 1: Connectivity.
 Role Prompt
@@
 The $k_{th}$ neighbor’s node index: $I_{k}^{j}$ Title: $\mathbf{t}_{title}^{I_{k}^{j}}$ Abstract: $\mathbf{t}_{abstract}^{I_{k}^{j}}$
-The full prompts for Connectivity is presented below.
+The full prompts for Connectivity is presented above.
 When generating prompts for different datasets, we adjust certain descriptions to better align with the specific context.
@@
 Output: The expanded training dataset $\tilde{\mathcal{D}_{L}}$, optimized parameters $LLM^{\theta^{\ast}},f_{S}^{\theta^{\ast}},f_{A}^{\theta^{\ast}}$, $f_{V}^{\phi^{\ast}}$ and predicted labels $\tilde{y}$.
-1 $\tilde{\mathcal{D}_{L}}\leftarrow\mathcal{D}_{L}$ // Training dataset initialization ;
-// LLM fine-tuning;
+1 $\tilde{\mathcal{D}_{L}}\leftarrow\mathcal{D}_{L}$;
 2 $LLM^{\theta^{\ast}}\leftarrow LLM^{\theta}$;
-// GNN-preference-driven Node Selector;
 3 Filter out GNN-preference nodes based on the preference rank $\mathcal{V}_{\mathcal{PR}}$ and get their annotations from $LLM^{\theta^{\ast}}$;
 4 Conduct prediction distillation from $LLM^{\theta^{\ast}}$ to $\{f_{T_{b}}^{\theta}\}_{b=1}^{4}$ for retrain them ;
-// Node-preference-driven GNN Selector;
 5 for $l_{1}\leftarrow 1$ to $L_{1}$ do
@@
 The specific analysis of the time complexity of PKD training and testing are provided below:
-The time complexity of PKD training is mainly divided into three parts: LLM fine-tuning, GNN-preference-driven Node Selector and Node-preference-driven GNN Selector.
+The time complexity of PKD training is mainly divided into three parts: LLM fine-tuning (Line 2), GNN-preference-driven Node Selector (Line 3-4) and Node-preference-driven GNN Selector (Line 5-12).
 The GNN-preference-driven Node Selector also can be divided into the annotations generation and prediction distillation.
 Table 12: The GTA fine-tuning configurations on Llama-3.1-8B-Instruct.
-Dataset size Epoch lora_r lora_alpha optimizer learning rate GTA fine-tuning time LLM
-53,617 2 4 4 AdamW $1e^{-4}$ 9h 41m 48s Llama-3.1-8B-Instruct
+Model Name Dataset Size Epoch lora_r lora_alpha Optimizer Learning Rate Time Cost
+Llama-3.1-8B-Instruct 53,617 2 4 4 AdamW [57] $1e^{-4}$ 9h 41m 48s
 First, we use Low-Rank Adaptation (LoRA) strategy [58] for efficient parameter training, with hyperparameters set to $r=4$, $\alpha=4$, $epoch=2$ (as shown in Table 12), and the rest are set according to the default settings of llama-factory [^1] https://github.com/hiyouga/LLaMA-Factory.
@@
 $r\ll d$, so the time complexity is bound by $\mathcal{O}(NLd+Ld^{2})$.
+Table 13: The time costs on Cora and Ogbn-Arxiv of annotations generation by Llama-3.1-8B-Instruct.
+Dataset Cora Ogbn-Arxiv
+Time / GPU-hours 0.11 1.68
 The process of annotations generation includes sorting the selected nodes and the reasoning process of LLM, and its time complexity is $\mathcal{O}(W\log W)$ and $\mathcal{O}(WL^{\prime}(l^{2}d+ld^{2}))$, where $W$ is the number of selected nodes, $L^{\prime}$ is the number of transformer layers in LLM, and $l$ is the input sequence length.
 Generally, $W\ll l$, $L^{\prime}\ll l$, then the time complexity is $\mathcal{O}(l^{2}d+ld^{2})$.
+Table 14: The time costs of retraining teacher GNNs on Cora and Ogbn-Arxiv.
+Datasets Teacher GNNs ($T_{1},T_{2},T_{3},T_{4}$) Total running time (seconds)
+Cora GCN, GAT, APPNP, H2GCN 2.4716
+Obgn-Arxiv GCN, GAT, APPNP, H2GCN 18.2017
 The time complexity of teacher GNN (2-layers) re-training is bound by $\mathcal{O}((NF+M)D)$, where $N$ is the number of nodes, $F$ is the node feature dimension, $M$ is the number of edges, and $D$ is the GNN hidden layer dimension.
+Table 15: Running times of Node-preference-driven GNN Selector on Cora and Obgn-Arixv. "m" and "s" denote minute and second.
+Dataset Peak memory Running time / epoch
+Cora 454.62 MB 7.9s
+Obgn-Arixv 1655.18 MB 44m 8.3s
 The time complexity of Node-preference-driven GNN Selector is $\mathcal{O}(W(l^{2}d+ld^{2}+dd’+d^{\prime}a))$, where $d^{\prime}$ is the dimension of the MLP hidden layer, $a$ is the number of action categories, $W\ll l$, $a\ll d$, so the time complexity is bound by $\mathcal{O}(l^{2}d+ld^{2}+dd^{\prime})$.
-The training time complexity of student GNN is $\mathcal{O}((NF+M)D)$ Therefore, the overall time complexity is bound by $\mathcal{O}((L+2l)d^{2}+(d’+nL)d+2l^{2}+2D(NF+M))$.
+The training time complexity of student GNN is $\mathcal{O}((NF+M)D)$.
+Therefore, the overall time complexity is bound by $\mathcal{O}((L+2l)d^{2}+(d’+nL)d+2l^{2}+2D(NF+M))$.
 The inference time complexity of PKD is determined by the testing process of the student GNN. So its time complexity is bound by $\mathcal{O}((NF+M)D)$.
@@
 The peak memories of performing PKD on the Cora and Ogbn-Arxiv are also listed in Table 15.
-Table 13: The time costs on Cora and Ogbn-Arxiv of different LLMs.
-Dataset Llama-3.1 Qwen-2.5
-Parameters 8B 13B 7B 14B
-Cora 0.11 GPU-hours 0.36 GPU-hours 0.05 GPU-hours 0.29 GPU-hours
-Ogbn-Arxiv 1.68 GPU-hours 5.01 GPU-hours 0.94 GPU-hours 2.95 GPU-hours
-Table 14: The time costs of retraining teacher GNNs on Cora, Amazon Ratings, and Ogbn-Arxiv.
-Datasets Teacher GNNs Epochs Total running time (seconds)
-Cora GCN, GAT, APPNP, H2GCN 300 2.4716
-Obgn-Arxiv GCN, GAT, APPNP, H2GCN 300 18.2017
-Amazon Ratings DirGNN, GPRGNN, HoloNets, H2GCN 300 3.1452
-Table 15: Running times of PPO on Cora and Obgn-Arixv. "m" and "s" denote minute and second.
-Dataset Peak memory Running time / epoch
-Cora 454.62 MB 7.9s
-Obgn-Arixv 1655.18 MB 44m 8.3s
 Appendix D Proofs for Propositions 3.1
@@
 Compared to these baselines, our method generates embeddings with significantly enhanced inter-class separability and high cluster purity, resulting in improved few-shot node classification performance.
-(a) Original student GNN
-(b) GCNII
-(c) EGNN
-(d) LLMGNN
-(e) GAugLLM
-(f) AGST
-(g) KDGA
-(h) MSKD
-(i) BGNN
-(j) MTAAM
-(k) FairGKD
-(l) PKD # LN 5
 Figure 6: T-SNE [63] visualizations on Cora.
 E.2 Ablation Study
-The results on Cora and Amazon Ratings of our RL-based method (PKD) and the other three teacher selection methods (entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, random selection, and end-to-end learning) are shown in Table 16.
+The results on Cora and Amazon Ratings using our PKDLlama (RL-based method) and the other three teacher selection methods (Entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, Random selection, and End-to-end learning) are shown in Table 16.
 It is obvious that the RL-based method significantly outperforms the other three methods.
@@
 End-to-end learning 60.29 60.39
-RL-based (ours) 90.27 65.93
+PKDLlama (RL-based) 90.27 65.93
 E.3 Hyperparameters Sensitivity Analysis
 As mentioned in Sec.
-4.4, we perform the hyperparameter sensitivity analysis over the loss-weight coefficients $\alpha,\beta,\gamma,\eta$ on two datasets, and report the results in Tables 17 18 19 20.
+4.4, we perform the hyperparameter sensitivity analysis over the loss-weight coefficients $\alpha,\beta,\gamma,\eta$ on two datasets, and report the results in Tables 171819 and 20.
 As result, the proposed PKD can achieve much better performance when $\alpha=0.5,\beta=1,\gamma=0.1,\eta=0.3$.
 Table 17: The influence of $\alpha$.
-$\alpha$ 2 1 0.9 0.8 0.7 0.6 0.5 0.4 0.3
-Cora 70.12 69.49 72.71 69.05 72.41 73.15 75.66 74.70 73.78
-Amazon Ratings 61.74 61.66 60.47 60.28 61.30 62.76 63.83 62.86 61.81
+$\alpha$ 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1 2
+Cora 73.78 74.70 75.66 73.15 72.41 69.05 72.71 69.49 70.12
+Amazon Ratings 61.81 62.86 63.83 62.76 61.30 60.28 60.47 61.66 61.74
 Table 18: The influence of $\beta$.
-$\beta$ 2 1 0.5 0.25
-Cora 74.15 75.36 67.17 64.10
-Amazon Ratings 63.93 63.91 63.23 62.94
+$\beta$ 0.25 0.5 1 2
+Cora 64.10 67.17 75.36 74.15
+Amazon Ratings 62.94 63.23 63.91 63.93
 Table 19: The influence of $\gamma$.
-$\gamma$ 2 1 0.5 0.2 0.1 0.05
-Cora 62.66 65.06 64.40 63.77 70.12 68.90
-Amazon Ratings 63.93 63.93 63.63 63.71 64.01 62.58
+$\gamma$ 0.05 0.1 0.2 0.5 1 2
+Cora 68.90 70.12 63.77 64.40 65.06 62.66
+Amazon Ratings 62.58 64.01 63.71 63.63 63.93 63.93
 Table 20: The influence of $\eta$.
-$\eta$ 0.9 0.7 0.5 0.3 0.1
-Cora 89.42 86.67 88.43 90.27 89.96
-Amazon Ratings 63.81 63.75 64.79 65.93 64.87
+$\eta$ 0.1 0.3 0.5 0.7 0.9
+Cora 89.96 90.27 88.43 86.67 89.42
+Amazon Ratings 64.87 65.93 64.79 63.75 63.81
 Appendix F Broader Impact